1.    Calculate the damping constant of an under damped harmonic oscillator, if its amplitude reduces to (1/10)th of its initial value after 100 oscillations. Given that time period is 2 s.



1
Expert's answer
2020-10-25T18:28:25-0400

The amplitude of a under damped harmonic oscillations is proportional to:


"A\propto \exp(-\omega\zeta t)"

where "\zeta" is the damping constant, "t" is a time and "\omega = \dfrac{2\pi}{2s} = \pi\space rad/s" is the angular frequency. If the periond of oscillation is 2s then 100 oscillations takes "\Delta t = 100\times 2s = 200s". In this time the amplitude reduces to (1/10)th of its initial value:


"\dfrac{A_2}{A_1} = \exp(-\omega\zeta (t_2-t_1)) = \exp(-\omega\zeta \Delta t) = \dfrac{1}{10}"

Expressing "\zeta", obtain:


"\zeta = \dfrac{\ln10}{\omega \Delta t} = \dfrac{\ln10}{\pi\cdot 200} \approx 0.0037"

Answer. 0.0037.